AUS Rural Bank is studying the relationship between the mean account balance for individual accounts and the number of transactions per month. A sample of eight accounts revealed:


customer 1 2 3 4 5 6 7 8


Mean balance 50 32 53 24 15 10 16 27


No of transactions 5 2 7 9 10 2 4 11



Find the coefficient of correlation. Determine at the 0.01 significance level whether the correlation in the population is greater than zero.



Step 1: state the hypotheses,



Ho: __________________________________________________________


H1: __________________________________________________________



Step 2: The level of significance and critical region: α=__________


And t critical______________


Step 3: Complete the table and compute for the value of r and t


customer X Y X2 Y2 XY


1


2


3


4


5


6


7


8


∑▒〖X=〗


∑▒Y=


∑▒〖X^2=〗


∑▒〖Y^2=〗


∑▒〖XY=〗


r=


tcomputed=


Step 4: Decision rule:


Step 5: Conclusion



1
Expert's answer
2021-06-07T13:16:27-0400


The coefficient of correlation between two variables can be obtained by:

"r=\frac{\sum (y-\bar{y}) \times (x-\bar{x})}{\sqrt{\sum (x-]bar{x})^2 \times \sum (y-\bar{y})^2}} \\

= \frac{13.25}{\sqrt{87.5 \times 1777.875}} \\

= 0.0336"

H0: ρ=0

H1: ρ>0

The level of significance and critical region: α=0.01

n=8

d.f.=n-1=7

"t= \frac{r\times \sqrt{n-2}}{\sqrt{1-r^2}} \\

= \frac{0.0336 \times \sqrt{8-2}}{\sqrt{1 -0.0336^2}} \\

= \frac{0.0822881}{0.999436} \\

= 0.082335"

The p-value can be obtained as:

The chart value of probability for t=0.082335 with d.f.=7 is 0.5315

So, p-value for a right tailed test is

p-value=1-0.5315 = 0.4685

Decision rule: We reject the H0 if the p-value of the test is less than α=0.01.

P-value > 0.01

So, we do not have enough evidence against H0 to reject it, so we fail to reject the H0.

We can conclude, that the correlation in the population is not greater than zero.